Given the following information for Company A, find the WACC. Assume the company’s tax rate is 21%

i.                Debt: 15,000 bonds selling for RM 1080, with a Yield To Maturity 5%

ii.              Common Stock: 500,000 shares outstanding, selling for RM 50 per share, the beta is 1.09

iii.            Preferred stock: 30,000 shares, paying fixed RM 5 dividend per share, selling at RM 60 per stock


The market risk - free rate is at 3.2% with 7% market risk premium.


1
Expert's answer
2022-01-10T09:55:21-0500

we calculate the debt:

"16000\times1080\times105=18 144 000"


"CARM=3.2+1.09(7-3.2)=7.342"


we calculate the equity

"500 000\times50=25 000 000"


 preferred shares:

"30 000\times60=1 800 000"


"18 144 000+25 000 000+1 800 000=44 944 000"


"WACC=7.342\frac{25 000 000}{44 944 000}+\frac{5}{60}\frac{1 800 000}{44 944 000}+5\frac{18 144 000}{44 944 000}\times(1-0,21)=5.68"